OpenStudy (anonymous):

If x2 = 10, what is the value of x? ±square root of 10 ±square root of 20 ±5 ±20

OpenStudy (anonymous):

5*2=10 x=5

OpenStudy (timothywhitaker):

you divid 10 by 2

OpenStudy (anonymous):

thankyou guyss!

OpenStudy (mathstudent55):

\(x^2 = 10\) \(x^2 - 10 = 0\) \((x + \sqrt {10})(x - \sqrt{10}) =0\) \(x = -\sqrt{10}\) or \(x =\sqrt{10} \) \(x = \pm \sqrt{10} \)

OpenStudy (anonymous):

You just have divide 10 by 2!